OpenStudy (anonymous):

Determine whether 2 is a zero of the polynomial P(x) = 3x3 – 6x2 – 3x – 18 by using the Remainder Theorem. Explain how the Remainder Theorem is useful in finding the zeros of a polynomial function.

OpenStudy (anonymous):

P(x) = 3x3 – 6x2 – 3x – 18 using the remainder theorem P(2) = 3(2)^3 – 6(2)^2 – 3(2) – 18 p(2)=-24 the remainder r=-24 if you divide 3x3 – 6x2 – 3x – 18 by x-2

OpenStudy (anonymous):

3x^2 -3 - ------------------ x-2/3x^3 – 6x^2 – 3x – 18 -(3x^3 -6x^2) ------------------- -3x-18 -(-3x+6) -------- -24=R